OpenStudy (anonymous):

Solve ^3 sqrt x-3 – 5 = 16

OpenStudy (anonymous):

\[\sqrt[3]{(x-3)}-5=16\] is this the question?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

\[\sqrt[3]{(x-3)}=21\] \[(x-3) = (21)^{3}\] \[x-3=9261\] \[x = 9264\]

OpenStudy (anonymous):

My choices are 52 67 19 5 ..... im so confused:(

OpenStudy (anonymous):

huh.... is the equation is seriously cube root?? Check my first reply.. is the question correct?